31st August (Monday)

  • Data Processing:
    • Drop of MC jobs at T1 sites on the week-end, solved this morning and restarting now. DataReconstruction at T1s.

  • T0

  • T1
    • RAL: Problem with pilot submission to ARC-CEs, working on a fix
    • CNAF OUTAGE downtime declared Friday evening for LHCb storage

27th August (Thursday)

  • Data Processing:
    • p-Ne collisions reconstruction started

  • T0
    • Nothing to report

  • T1
    • RAL 'at-risk' yesterday and today due to castor update.

24th August (Monday)

  • Data Processing:
    • validation productions of 25ns data will be redone this week (ReCo + Stripping), with proper refitting of PV
    • current validation (Reco15b and Stripping 23a) quasi finished (>99% data processed, 75% merged)

  • T0
    • Nothing to report

  • T1
    • Nothing to report

20th August (Thursday)

  • Data Processing: second run of validation for 25ns data started with new calibration, ~90% already processed. New stripping production will start later today.

  • T0
    • Nothing to report

  • T1
    • Nothing to report
    • CVMFS problem at NIKHEF-ELPROD (GGUS:115767) due to an upgrade of their WNs from CentOS 6.6 to CentOS 6.7. Promptly solved.
    • PBS issues at PIC (GGUS:115748) due to a high load in few WNs, pbs server had problems to refresh information, and denied some connections. Promptly solved.

17th August (Monday)

  • Data Processing: validation of 25ns data started. ~70% already processed.

  • T0
    • Nothing to report

  • T1
    • Nothing to report

13th August (Thursday)

  • Data Processing: data "stripping" productions ongoing almost complete. Ready to process new data. User and MC jobs

  • T0
    • Nothing to report

  • T1
    • GRIDKA: Still we have problems getting TURLs from SURLs.

10th August (Monday)

  • Data Processing: data "stripping" productions ongoing almost complete. Ready to process new data. User and MC jobs

  • T0
    • Nothing to report

  • T1
    • RAL: Network issues over the weekend
    • GRIDKA: Problems over the weekend getting TURLs from SURLs. Overloaded SRM? Difficult to track down as the jobs retry/go to failover and complete.

6th August (Thursday)

  • Data Processing: data "stripping" productions ongoing. User and MC jobs

  • T0
    • Discussion with LSF team about wall and cpu time queue lengths, ongoing (GGUS:115027). Some progress and bugs found on both sides. Deployment of fixes happening in progress.

  • T1
    • RAL: All network problems appear to be solved.

3rd August (Monday)

  • Data Processing: data "stripping" productions ongoing. User and MC jobs

  • T0
    • Discussion with LSF team about wall and cpu time queue lengths, ongoing (GGUS:115027). Some progress and bugs found on both sides. Deployment of fixes happening in progress.

  • T1
    • RAL: LAN network instabilities continue, CVMFS and data access problems over the weekend.

-- JoelClosier - 2016-01-07

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2016-01-07 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback